OpenStudy (anonymous):

.060 ball moving @ 2.5 m/s to the right collides with a .090 ball @ rest. How are initial and final momenta related? What are speeds and directions of each after collision? Impulse imparted during collision?

OpenStudy (anonymous):

pi=.060*2.5= pf=(.090+.060)*vf pf=pi => vf= I=(.090+.060)*vf-0.60*2.5= (not sure)